OpenStudy (anonymous):

the entire surface of a cube is 1014 sq ft find the volume and the length of one diagonal please help!

OpenStudy (anonymous):

6x^2 = 1014

OpenStudy (amistre64):

1014/6 gives you the area of one face

OpenStudy (amistre64):

sqrt(face) = sedge

OpenStudy (dumbcow):

diagonal = 13sqrt(2) :)

OpenStudy (amistre64):

sqrt(3edge^2) = diagonal from corner to corner

OpenStudy (anonymous):

\[\sqrt{169} =13\]

OpenStudy (dumbcow):

volume = 13^3

OpenStudy (anonymous):

13^3 will give you the volume

OpenStudy (anonymous):

ThaNk youu :)